OpenStudy (winsteria):

Which equation is related to the equation shown? 5x = 35 A. x = 35+ 5 B. x = 35÷5 C. x = 35-5 D. x = 35 x 5 CAN SOMEONE PLEASE HELP ME I'LL GIVE YOU MEDAL AND FAN

OpenStudy (winsteria):

CAN SOMEONE PLEASE HELP ME

OpenStudy (kkutie7):

b

OpenStudy (kkutie7):

do to one side what you do to the other. on the left we want x alone so dividing by 5 gets rid of 5. do the same to the other side.

OpenStudy (winsteria):

okay thank you can you help me with another one?

OpenStudy (kkutie7):

sure

OpenStudy (winsteria):

Which equation represents the word sentence? The product of 6 and a number increased by 4 is 28. A. 6x + 4 = 28 B. 6x + 4 + 28 C. 6x – 4 = 28 D. 6x + 28 = 4

OpenStudy (kkutie7):

a

OpenStudy (kkutie7):

a number is x and product is multiplication. when increasing you add.

OpenStudy (winsteria):

thank you so much

OpenStudy (winsteria):

can you help me with this one too?

OpenStudy (winsteria):

What is the value of the expression when a = 6 and b = 4? 14a - 2(3b + 4) please???

OpenStudy (anonymous):

1312